How many ways are there to seat 8 people, consisting of 4 couples, in a row of seats (8 seats wide) if all couples are to get adjacent seats?
Answer #1

4 couples can sit in 4! = 4 * 3 * 2 * 1 = 24 ways.

Now in each couple 2 people can sit in 2! = 2 * 1 = 2 ways.

So total number of ways = 24 * 2 = 48 (Ans)

** If the answer does not match please comment.